OpenStudy (anonymous):

Which expression is the simplified form of the square root of c to the fifteenth power.? c3the square root of c to the fifth power. c7the square root of c. c5the square root of c to the fifth power. c3the square root of c to the sixth power.

OpenStudy (anonymous):

\[(\sqrt{c})^{15}=c^{\frac{15}{2}}=c^{7+\frac{1}{2}}=c^7*c^{\frac{1}{2}}=c^7*\sqrt{c}\]

OpenStudy (anonymous):

thats ur answer written mathematically so u just have to find which answer is the same as the 1 i wrote